The least of 3 consecutive integers is divided by 10, the next is divided by 17, and the greatest is divided by 26. What are the numbers if the sum of the quotient is 10?
Answer by stanbon(48558)

The least of 3 consecutive integers is divided by 10, the next is divided by 17, and the greatest is divided by 26. What are the numbers if the sum of the quotient is 10?
--------------------
1st: x-1
2nd: x
3rd: x+1
---------------
EQUATION:
(x-1)/10 + x/17 + (x+1)/26 = 10
17*26(x-1) + 10*26x + 10*17(x+1) + 10*10*17*26
872x -272 = 44200
872x = 44472
x = 51 (2nd)
x-1= 50 (1st)
x+1 = 52 (3rd)
====================
Cheers,
Stan H.